# Extremely Amenable Automorphism Groups of Countable Structures

### v1

@@ 3.1. $\Delta$-metric spaces
Let $\Delta$ be a set of positive real numbers. We call $\Delta$ a distance value set if for all $x,y\in\Delta$,
@@
Fix a countable distance value set $\Delta$. We can view each countable $\Delta$-metric space as a countable structure, as follows. Let $\mathcal{L}_{\Delta}=\{R_{s}\colon s\in\Delta\}$, where each $R_{s}$ is a binary relation symbol. Given any countable $\Delta$-metric space $(X,d)$, set $R^{X}_{s}(x,y)$ iff $d(x,y)=s$. Thus we turn $(X,d)$ into a countable $\mathcal{L}_{\Delta}$-structure $(X;(R^{X}_{s})_{s\in\Delta})$. Let $\mbox{\rm Aut}(X,d)$, or simply $\mbox{\rm Aut}(X)$ if $d$ is understood, denote the automorphism group $\mbox{\rm Aut}(X;(R^{X}_{s})_{s\in\Delta})$. Then $\mbox{\rm Aut}(X)$ is a Polish group.
@@
Importantly, note that $\mbox{\rm Aut}(X)$ and $\mbox{\rm Iso}(X)$ are isomorphic as abstract groups.
@@ 3.2. The extreme amenability of $\mbox{\rm Aut}(\mathbb{U}_{\Delta}^{<})$
To apply the KPT correspondence, we need to consider a Fraïssé ordered class. Let $\Delta$ be a countable distance value set. Let $\mathcal{K}_{\Delta}^{<}$ be the class of all finite ordered $\Delta$-metric spaces, i.e., each element of $\mathcal{K}_{\Delta}^{<}$ is a structure $(X;(R_{s}^{X})_{s\in\Delta},<^{X})$, where $<^{X}$ is a linear order on $X$. It is easy to see that $\mathcal{K}_{\Delta}^{<}$ is a Fraïssé (ordered) class. We denote its Fraïssé limit by $\mathbb{U}_{\Delta}^{<}$. It can be shown that $\mathbb{U}_{\Delta}^{<}\cong(\mathbb{U}_{\Delta},<^{\mathbb{U}_{\Delta}})$ where $<^{\mathbb{U}_{\Delta}}$ is a linear order isomorphic to $(\mathbb{Q},<)$ (see, e.g., Proposition 5.2 of [15]).
@@ 3.3. The extreme amenability of $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$
Note that every automorphism of $\mathbb{U}_{\Delta}^{<}$ is actually an isomertry of $\mathbb{U}_{\Delta}$. So we consider the identity map from ${\rm Aut}(\mathbb{U}_{\Delta}^{<})$ to ${\rm Iso}(\mathbb{U}_{\Delta})$. This map is obviously continuous. The following proposition implies that $\mbox{\rm Aut}(\mathbb{U}^{<}_{\Delta})$ is dense in $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$.
@@
Suppose $\Delta$ is a countable distance value set with $\inf\Delta=0$. Given $x_{1},\dots,x_{n},y_{1},\dots,y_{n}\in\mathbb{U}_{\Delta}$ such that $x_{i}\mapsto y_{i}$, $1\leq i\leq n$, is a partial isometry, and given $\epsilon>0$, there exist $y_{1}^{\prime},\dots,y_{n}^{\prime}\in\mathbb{U}_{\Delta}$ so that $x_{i}\mapsto y_{i}^{\prime}$, $1\leq i\leq n$, is an order-preserving (with respect to $<^{\mathbb{U}_{\Delta}}$) partial isometry, and for all $1\leq i\leq n$, $d(y_{i},y_{i}^{\prime})<\epsilon$.
@@
    ${\rm Iso}(\mathbb{U}_{\Delta})$ is extremely amenable iff $\inf\Delta=0$.
@@
    ${\rm Iso}(\mathbb{U}_{\Delta})$ is Polish iff $\inf\Delta>0$.
@@
For (1) first assume that $\inf\Delta=0$. Then from Proposition 3.3, we conclude that $\mbox{\rm Aut}(\mathbb{U}_{\Delta}^{<})$ is dense in $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$. Now by Theorems 3.1 and 2.1, $\mbox{\rm Aut}(\mathbb{U}^{<}_{\Delta})$ is extremely amenable. Thus $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ is extremely amenable by Lemma 3.2.
@@
Next we prove (3). It was proved in [3] that if $\Delta=\mathbb{Q}$, then $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ is dense in $\mbox{\rm Iso}(\mathbb{U})$. The argument in [3] can be generalized to prove that if $\Delta$ is dense in $(0,+\infty)$ then ${\rm Iso}(\mathbb{U}_{\Delta})$ is a dense subgroup of ${\rm Iso}(\mathbb{U})$. Then same argument also shows that when $\inf\Delta=0$ and $\Delta$ is bounded, $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ is a dense subgroup of $\mbox{\rm Iso}(\mathbb{U}_{a})$, where $a=\sup\Delta$ and $\mathbb{U}_{a}$ is a sphere of $\mathbb{U}$ with diameter $a$. However, for any $0<a<\infty$, $\mbox{\rm Iso}(\mathbb{U}_{a})$ and $\mbox{\rm Iso}(\mathbb{U}_{1})$ are isomorphic. Hence in this case $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ is isomorphic to a dense subgroup of $\mbox{\rm Iso}(\mathbb{U}_{1})$.
@@
By Theorem 3.4 (1) and (2), for any countable distance value set $\Delta$, the separable metrizable group ${\rm Iso}(\mathbb{U}_{\Delta})$ can never be both Polish and extremely amenable. When $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ is not Polish, one typically takes the metric completion of $\mathbb{U}_{\Delta}$ and considers its isometry group, which is Polish. Vershik [32] (Theorem 3) has shown that if $\Delta$ is dense in $(0,+\infty)$, then $\mathbb{U}_{\Delta}$ is dense in $\mathbb{U}$, hence the metric completion of $\mathbb{U}_{\Delta}$ is $\mathbb{U}$. Similarly, if $\Delta$ is bounded and dense in $(0,\sup\Delta)$, then the metric completion of $\mathbb{U}_{\Delta}$ is just $\mathbb{U}_{a}$ where $a=\sup\Delta$. Thus, clause (3) of Theorem 3.4 says that the above-mentioned consideration obtains essentially only two Polish groups, $\mbox{\rm Iso}(\mathbb{U})$ and $\mbox{\rm Iso}(\mathbb{U}_{1})$, both known to be extremely amenable by Pestov [25].
4. Equivalence of Distance Value Sets
In this section, we deviate from the extremely amenable groups and study a notion of equivalence between distance value sets.
@@ 4. Equivalence of Distance Value Sets
    We say that $\Delta$ and $\Lambda$ are equivalent if there is a bijection $f\colon\Delta\to\Lambda$ such that for any positive real numbers $x,y,z$, $(x,y,z)$ is a $\Delta$-triangle iff $(f(x),f(y),f(z))$ is a $\Lambda$-triangle.
Let $\Delta$ and $\Lambda$ be equivalent distance value sets. Assume that $\inf\Delta=0$. Suppose $f:\Delta\rightarrow\Lambda$ witnesses that they are equivalent. Fix a sequence $(\delta_{i})_{i\in\mathbb{N}}$ in $\Delta$ such that $\delta_{i+1}<{\delta_{i}}/{2}$ for all $i$.
Lemma 4.1.
$f(\delta_{i+1})<{f(\delta_{i})}/{2}$ for all $i$. Consequently, $\inf\Lambda=0$.
Proof.
For each $i$, consider the triple $(\delta_{i+1},\delta_{i+1},\delta_{i})$. This triple is not a $\Delta$-triangle since $\delta_{i}>\delta_{i+1}+\delta_{i+1}$. By the property of $f$, the triple $(f(\delta_{i+1}),f(\delta_{i+1}),f(\delta_{i}))$ is not a $\Lambda$-triangle. But then the only failed triangle inequality must be $f(\delta_{i})>f(\delta_{i+1})+f(\delta_{i+1})$. ∎
Lemma 4.2.
$f$ is continuous.
Proof.
Fix $x\in\Delta$. For any $\epsilon>0$, find an $i$ such that $\delta_{i}<x$ and $f(\delta_{i})<\epsilon$. For any $y\in\Delta$ with $|x-y|<\delta_{i}$, $(x,y,\delta_{i})$ is a $\Delta$-triangle. Thus $(f(x),f(y),f(\delta_{i}))$ is a $\Lambda$-triangle. In particular, $|f(x)-f(y)|<f(\delta_{i})<\epsilon$. ∎
The above lemma actually shows that for all positive $\delta$, $f$ is uniformly continuous on $\Delta^{>\delta}\triangleq\Delta\cap(\delta,+\infty)$. The following lemma shows that $f$ is in fact linear. In the proof we use notation such as $\Delta^{>\delta}$, $\Delta^{<\delta}$, and $\Delta^{[a,b]}$ with obvious meanings.
Lemma 4.3.
If $x,y,x+y\in\Delta$, then $f(x+y)=f(x)+f(y)$.
Proof.
Fix such $x,y\in\Delta$. We claim that
  1. (a)
    $f(\Delta^{[|x-y|,x+y]})\subseteq\Lambda^{[|f(x)-f(y)|,f(x)+f(y)]}$,
  2. (b)
    $f(\Delta^{>x+y})\subseteq\Lambda^{>f(x)+f(y)}\mbox{ or }f(\Delta^{>x+y})\subseteq\Lambda^{<|f(x)-f(y)|}$,
  3. (c)
    $f(\Delta^{<|x-y|})\subseteq\Lambda^{>f(x)+f(y)}\mbox{ or }f(\Delta^{<|x-y|})\subseteq\Lambda^{<|f(x)-f(y)|}$.
Let $z\in\Delta$. (a) is equivalent to saying that if $(x,y,z)$ is a $\Delta$-triangle, then $(f(x),f(y),f(z))$ is a $\Lambda$-triangle. If $z>x+y$, then $(x,y,z)$ is not a $\Delta$-triangle. Thus $(f(x),f(y),f(z))$ is not a $\Lambda$-triangle, and so either $f(z)>f(x)+f(y)$ or $f(z)<|f(x)-f(y)|$. Similarly for $z<|x-y|$.
To prove (b), assume $z_{1},z_{2}>x+y$ such that $f(z_{1})<|f(x)-f(y)|$ and $f(z_{2})>f(x)+f(y)$. Without loss of generality assume $z_{1}<z_{2}$ (the argument for the case $z_{1}>z_{2}$ is similar). By the continuity of $f$ and the density of both $\Delta$ and $\Lambda$, we may find $z_{1}=a_{1},\dots,a_{n}=z_{2}$ in $\Delta$ such that $a_{k}>x+y$ and $|f(a_{k+1})-f(a_{k})|<f(x)+f(y)-|f(x)-f(y)|$ for all $1\leq k\leq n$. Then there must be an $a_{k}$ with $|f(x)-f(y)|\leq f(a_{k})\leq f(x)+f(y)$, contradicting $a_{k}>x+y$. This proves (b).
A similar argument for $z_{1},z_{2}<|x-y|$ proves (c).
Since $f$ is a bijection from $\Delta$ to $\Lambda$, the claimed inclusions are actually equalities. By considering sufficiently small $\delta_{i}$, we see that $f(\Delta^{<|x-y|})=\Lambda^{<|f(x)-f(y)|}$, and thus $f(\Delta^{>x+y})=\Lambda^{>f(x)+f(y)}$. By the continuty of $f$ again, we have $f(x+y)=f(x)+f(y)$. ∎
Lemma 4.4.
Let $x,t\in\Delta$. We have $f(x)/x=f(t)/t$.
Proof.
For each $k>0$, find $0<\delta_{k}<1/2^{k}$ by the continuity of $f$ such that for all $z\in\Delta$, $|t-z|<\delta_{k}$ implies $|f(t)-f(z)|<1/k$. Now for each $k$, find positive rational number $n_{k}/m_{k}$ such that $20<t-(n_{k}/m_{k})x<\delta_{k}/$ 2. Note that by the density of $\Delta$ and bijectivity of $f$, $\delta_{k}\to 0$ as $k\to\infty$. Thus $n_{k}/m_{k}\to t/x$ as $k\to\infty$.
Letting $k\to\infty$, we get $f(t)=(t/x)f(x)$ and thus $f(x)/x=f(t)/t$. ∎
Thus we have shown the following theorem.
Theorem 4.5.
Let $\Delta$ and $\Lambda$ be distance value sets with $\inf\Delta=\inf\Lambda=0$. Then the following are equivalent:
  1. (1)
    $\Delta$ and $\Lambda$ are equivalent.
  2. (2)
    There is a positive real number $\alpha$ such that $\Lambda=\alpha\cdot\Delta$.
We consider a special kind of countable distance value sets for which the equivalence is more explicit.
Definition 4.2.
Let $\alpha$ be a positve irrational number. Let
$\Delta_{\alpha}=\{p\alpha+q:p,q\in\mathbb{Q}\}\cap(0,+\infty).$
Define an action of $GL_{2}(\mathbb{Q})$ on the space of irrational numbers by the linear fractional transfomation:
Theorem 4.6.
  1. (1)
    $\Delta_{\alpha}$ and $\Delta_{\beta}$ are equivalent.
  2. (2)
    $\alpha$ and $\beta$ are $E_{GL_{2}(\mathbb{Q})}^{IR}$-equivalent.
Proof.
Corollary 4.7.
There are continuum many positive irrational numbers which are pairwise $E_{GL_{2}(\mathbb{Q})}^{IR}$-inequivalent. Consequently, there are continnum many pairwise inequivalent countable distance value sets each without a positive lower bound.
Proof.
Note that each $E_{GL_{2}(\mathbb{Q})}^{IR}$-equivalence class is countable. Thus there are continuum many pairwise $E_{GL_{2}(\mathbb{Q})}^{IR}$-inequivalent positive irrational numbers. ∎
5. Isomorphism Types of Extremely Amenable Groups
@@ 5.1. Isomorphism types of $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$
In this subsection we show that there are continuum many pairwise non-isomorphic extremely amenable groups that are separable metrizable. The class of such groups are of the form $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ for countable distance value sets $\Delta$ with $\inf\Delta=0$. We use the following theorem from [4].
@@
Corollary 5.3.
There are continuum many pairwise non-isomorphic extremely amenable groups which are separable metrizable but not Polish.
Proof.
This follows immediately from Theorem 5.2, Corollary 4.7, and Theorem 3.4. ∎
5.2. Isomorphism types of $\mbox{\rm Aut}(\mathbb{U}^{<}_{\Delta})$
In this subsection we show that there are continuum many pairwise non-isomorphic extremely amenable Polish groups. This is done by considering $\mbox{\rm Aut}(\mathbb{U}^{<}_{\Delta})$ for countable distance value sets $\Delta$. Our main theorem of this subsection is the following.
@@ 5.2. Isomorphism types of $\mbox{\rm Aut}(\mathbb{U}^{<}_{\Delta})$
Granting the theorem, we immediately get the following corollary from Corollary 4.7.
Corollary 5.5.
There are continuum many pairwise non-isomorphic extremely amenable Polish groups which are automorphism groups of countable structures, or equivalently closed subgroups of $S_{\infty}$.
@@
Let $D$ be a finite ordered metric space. Let $p_{1}$ and $p_{2}$ be order-preserving partial isometries of $D$ such that $Z(p_{1})=F(p_{1})$ and $Z(p_{2})=F(p_{2})$. Suppose $F(p_{1})\cap F(p_{2})\neq\emptyset$. Then there exist a finite ordered metric space $D^{\prime}$ extending $D$, order-preserving partial isometries $q_{1}$ and $q_{2}$ of $D^{\prime}$ extending $p_{1},p_{2}$ respectively, and an element $w\in\mathsf{F}(s,t)$, the free group generated by ${s,t}$, such that
@@
Intuitively, $g_{1}$ sends $w(y)$ to a copy which is smaller than $b_{n}$ and fixes $A$ pointwise. Consequently, we have $w_{1}(D)<b_{n}$ and, for any $y\in D$, $w_{1}(y)<x_{0}$ iff $w(y)<x_{0}$.
@@
Define an order-preserving partial isometry $g_{2}\colon B\cup w_{1}(D)\to B\cup w_{2}(D)$ by letting $g_{2}(b)=b$ for all $b\in B$ and $g_{2}(w_{1}(y))=w_{2}(y)$ for all $y\in D$. Moreover, $w_{2}(D)<a_{n}$ and, for any $y\in D$, $w_{2}(y)<x_{0}$ iff $w_{1}(y)<x_{0}$.
@@
We also need the following two lemmas about proper open subgroups of $\mbox{\rm Aut}(\mathbb{U}^{<}_{\Delta})$.
@@
Toward a contradiction, assume $V\cdot x$ is infinite for all $x\in X$. We prove that $V$ is dense. Since $V$ is also closed, it follows that $V={\rm Aut}(\mathbb{U}_{\Delta}^{<})$, contradicting the properness of $V$.
Note that a basic open neighborhood of the identity element of $\mbox{\rm Aut}(\mathbb{U}^{<}_{\Delta})$ has the form ${\rm Aut}_{A}(\mathbb{U}_{\Delta}^{<})$, where $A$ is a finite subset of $\mathbb{U}^{<}_{\Delta}$. Since $V$ is an open subgroup, there is a finite $A\subseteq\mathbb{U}^{<}_{\Delta}$ such that ${\rm Aut}_{A}(\mathbb{U}_{\Delta}^{<})\subseteq V$. By Neumann’s lemma (see Corollary 4.2.2 of Hodges [13]), there exists $g\in V$ such that $g(A)\cap A=\varnothing$.
@@
First suppose $V={\rm Aut}_{\{x\}}(\mathbb{U}_{\Delta}^{<})$. Assume $H$ is a closed subgroup of $\mbox{\rm Aut}(\mathbb{U}^{<}_{\Delta})$ such that $V\subsetneq H\subsetneq\mbox{\rm Aut}(\mathbb{U}^{<}_{\Delta})$. Let $g\in H\setminus V$. Then $g(x)\neq x$ and ${\rm Aut}_{\{g(x)\}}(\mathbb{U}_{\Delta}^{<})=g{\rm Aut}_{\{x\}}(\mathbb{U}_{\Delta}^{<})g^{-1}=gVg^{-1}\leq H$. By Lemma 5.8, $\langle{\rm Aut}_{\{g(x)\}}(\mathbb{U}_{\Delta}^{<}),{\rm Aut}_{\{x\}}(\mathbb{U}_{\Delta}^{<})\rangle$ is dense in ${\rm Aut}(\mathbb{U}_{\Delta}^{<})$. Since it is a subgroup of $H$ and $H$ is closed, we have $H={\rm Aut}(\mathbb{U}_{\Delta}^{<})$, contradicting the properness of $H$.
For the converse, assume (2) holds. By Lemma 5.9, there exists an element $x\in\mathbb{U}_{\Delta}^{<}$ such that $V\cdot x$ is finite. We show that $V\subseteq\mbox{\rm Aut}_{\{x\}}(\mathbb{U}^{<}_{\Delta})$. Otherwise, let $g\in V$ be such that $g(x)\neq x$. Then either $g(x)>x$ or $g(x)<x$. Since $g$ is order-preserving, we get that either $g^{n+1}(x)>g^{n}(x)$ for all natural numbers $n$, or $g^{n+1}(x)<g^{n}(x)$ for all natural numbers $n$. In either case, $V\cdot x$ is infinite.
It is clear that $\mbox{\rm Aut}_{\{x\}}(\mathbb{U}^{<}_{\Delta})\neq\mbox{\rm Aut}(\mathbb{U}^{<}_{\Delta})$. Thus by the the maximality of $V$, we have $V=\mbox{\rm Aut}_{\{x\}}(\mathbb{U}^{<}_{\Delta})$.
@@
Let $f\colon\mbox{\rm Aut}(\mathbb{U}^{<}_{\Delta})\to\mbox{\rm Aut}(\mathbb{U}^{<}_{\Lambda})$ be a topological group isomorphism. By Lemma 5.10, for each $x\in\mathbb{U}_{\Delta}^{<}$, there is a unique $y\in\mathbb{U}_{\Lambda}^{<}$ such that $f({\rm Aut}_{\{x\}}(\mathbb{U}_{\Delta}^{<}))={\rm Aut}_{\{y\}}(\mathbb{U}_{\Lambda}^{<})$. We write $y=g(x)$. The map $g$ is well-defined and bijective.
Note that for any $x,y,x^{\prime},y^{\prime}\in\mathbb{U}_{\Delta}^{<}$, ${\rm Aut}_{\{x,y\}}(\mathbb{U}_{\Delta}^{<})$ and ${\rm Aut}_{\{x^{\prime},y^{\prime}\}}(\mathbb{U}_{\Delta}^{<})$ are conjugate in $\mbox{\rm Aut}(\mathbb{U}^{<}_{\Delta})$ iff $d(x,y)=d(x^{\prime},y^{\prime})$. Also note that
${\rm Aut}_{\{x,y\}}(\mathbb{U}_{\Delta}^{<})=\mbox{\rm Aut}_{\{x\}}(\mathbb{U}^{<}_{\Delta})\cap\mbox{\rm Aut}_{\{y\}}(\mathbb{U}^{<}_{\Delta})$
and similarly for $\mbox{\rm Aut}_{\{x^{\prime},y^{\prime}\}}(\mathbb{U}^{<}_{\Delta})$. Thus by the definition of $g$, we have that for any $x,y,x^{\prime},y^{\prime}\in\mathbb{U}_{\Delta}^{<}$,

### v2

@@ 3.1. $\Delta$-metric spaces
Let $\Delta$ be a nonempty set of positive real numbers. We call $\Delta$ a distance value set if for all $x,y\in\Delta$,
@@
Fix a countable distance value set $\Delta$. We can view each countable $\Delta$-metric space as a countable structure, as follows. Let $\mathcal{L}_{\Delta}=\{R_{s}\colon s\in\Delta\}$, where each $R_{s}$ is a binary relation symbol. Given any countable $\Delta$-metric space $(X,d)$, set $R^{X}_{s}(x,y)$ if and only if $d(x,y)=s$. Thus we turn $(X,d)$ into a countable $\mathcal{L}_{\Delta}$-structure $(X;(R^{X}_{s})_{s\in\Delta})$. Let $\mbox{\rm Aut}(X,d)$, or simply $\mbox{\rm Aut}(X)$ if $d$ is understood, denote the automorphism group $\mbox{\rm Aut}(X;(R^{X}_{s})_{s\in\Delta})$. Then $\mbox{\rm Aut}(X)$ is a Polish group.
@@
Importantly, note that $\mbox{\rm Aut}(X)$ and $\mbox{\rm Iso}(X)$ are identical as sets, and the identity map is a group isomorphism.
@@ 3.2. The extreme amenability of $\mbox{\rm Aut}(\mathbb{U}_{\Delta}^{<})$
To apply the KPT correspondence, we need to consider a Fraïssé ordered class. Let $\Delta$ be a countable distance value set. Let $\mathcal{K}_{\Delta}^{<}$ be the class of all finite ordered $\Delta$-metric spaces, i.e., each element of $\mathcal{K}_{\Delta}^{<}$ is a structure $(X;(R_{s}^{X})_{s\in\Delta},<^{X})$, where $^{X}$ is a linear order on $X$. It is easy to see that $\mathcal{K}_{\Delta}^{<}$ is a Fraïssé (ordered) class. We denote its Fraïssé limit by $\mathbb{U}_{\Delta}^{<}$. It can be shown that $\mathbb{U}_{\Delta}^{<}\cong(\mathbb{U}_{\Delta},<^{\mathbb{U}_{\Delta}})$ where $^{\mathbb{U}_{\Delta}}$ is a linear order isomorphic to $(\mathbb{Q},<)$ (see, e.g., Proposition 5.2 of [17]).
@@ 3.3. The extreme amenability of $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$
[…]
@@
Suppose $\Delta$ is a countable distance value set with $\inf\Delta=0$. Given $x_{1},\dots,x_{n},y_{1},\dots,y_{n}\in\mathbb{U}_{\Delta}$ such that $x_{i}\mapsto y_{i}$, $1\leq i\leq n$, is a partial isometry, and given $\epsilon>0$, there exist $y_{1}^{\prime},\dots,y_{n}^{\prime}\in\mathbb{U}_{\Delta}$ so that $x_{i}\mapsto y_{i}^{\prime}$, $1\leq i\leq n$, is an order-preserving (with respect to $^{\mathbb{U}_{\Delta}}$) partial isometry, and for all $1\leq i\leq n$, $d(y_{i},y_{i}^{\prime})<\epsilon$.
@@
    ${\rm Iso}(\mathbb{U}_{\Delta})$ is extremely amenable if and only if $\inf\Delta=0$.
@@
    ${\rm Iso}(\mathbb{U}_{\Delta})$ is Polish if and only if $\inf\Delta>0$.
@@
[…]
@@
Next we prove (3). Recall that Vershik [36] (Theorem 3) has shown that if $\Delta$ is dense in $(0,+\infty)$, then $\mathbb{U}_{\Delta}$ is dense in $\mathbb{U}$, hence the metric completion of $\mathbb{U}_{\Delta}$ is $\mathbb{U}$. Similarly, if $\Delta$ is bounded and dense in $(0,\sup\Delta)$, then the metric completion of $\mathbb{U}_{\Delta}$ is just $\mathbb{U}_{a}$ where $a=\sup\Delta$ and $\mathbb{U}_{a}$ is a sphere of $\mathbb{U}$ with diameter $a$. Thus, given any isometry $f\in\mbox{\rm Iso}(\mathbb{U}_{\Delta})$, there is a unique isometry $\hat{f}\in\mbox{\rm Iso}(\mathbb{U})$ (or $\mbox{\rm Iso}(\mathbb{U}_{a})$ if $a=\sup\Delta<+\infty$) extending $f$. Therefore we can view $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ as the subgroup of $\mbox{\rm Iso}(\mathbb{U})$ (or of $\mbox{\rm Iso}(\mathbb{U}_{a})$) consisting of exactly those isometries fixing $\mathbb{U}_{\Delta}$ setwise. It was proved in [3] that if $\Delta=\mathbb{Q}$, then $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ is dense in $\mbox{\rm Iso}(\mathbb{U})$. The argument in [3] can be generalized to prove that if $\Delta$ is dense in $(0,+\infty)$ then ${\rm Iso}(\mathbb{U}_{\Delta})$ is a dense subgroup of ${\rm Iso}(\mathbb{U})$. The same argument also shows that when $\inf\Delta=0$ and $\Delta$ is bounded, $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ is a dense subgroup of $\mbox{\rm Iso}(\mathbb{U}_{a})$, where $a=\sup\Delta$. However, for any $0<a<\infty$, $\mbox{\rm Iso}(\mathbb{U}_{a})$ and $\mbox{\rm Iso}(\mathbb{U}_{1})$ are isomorphic. Hence in this case $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ is isomorphic to a dense subgroup of $\mbox{\rm Iso}(\mathbb{U}_{1})$.
@@
By Theorem 3.4 (1) and (2), for any countable distance value set $\Delta$, the separable metrizable group ${\rm Iso}(\mathbb{U}_{\Delta})$ can never be both Polish and extremely amenable. When $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ is not Polish, one typically takes the metric completion of $\mathbb{U}_{\Delta}$ and considers its isometry group, which is Polish. Clause (3) of Theorem 3.4 says that this obtains essentially only two Polish groups, $\mbox{\rm Iso}(\mathbb{U})$ and $\mbox{\rm Iso}(\mathbb{U}_{1})$, both known to be extremely amenable by Pestov [29].
We remark here that Theorem 3.4 (1) follows from a less well-known result of Gould. In [11, Theorem A.2.2 (4)] he showed that a topological group is extremely amenable if and only if all of its dense subgroups are extremely amenable. By this result, $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ is extremely amenable when $\inf\Delta=0$ following the density observations and results of Pestov mentioned in the preceding paragraph. The authors thank Yañez for pointing out the reference [11] to us.
In addition, Dai–Gao–Yañez [4] recently proved that for countable distance value sets $\Delta$ with $\inf\Delta=0$, $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ has the so-called strong Lévy property. This also implies that they are extremely amenable by a result of Gromov–Milman [12].
4. Isomorphism Types of Extremely Amenable Groups
[…]
@@ 4. Isomorphism Types of Extremely Amenable Groups
    We say that $\Delta$ and $\Lambda$ are equivalent if there is a bijection $f\colon\Delta\to\Lambda$ such that for any positive real numbers $x,y,z$, $(x,y,z)$ is a $\Delta$-triangle if and only if $(f(x),f(y),f(z))$ is a $\Lambda$-triangle.
@@ 4.1. Isomorphism types of $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$
In this subsection we characterize the topological isomorphism relation of groups of the form $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ for countable distance value sets $\Delta$ with $\inf\Delta=0$. We use the following theorem from [5].
@@
[…]
@@
Let $D$ be a finite ordered metric space. Let $p_{1}$ and $p_{2}$ be order-preserving partial isometries of $D$ such that $Z(p_{1})=F(p_{1})$ and $Z(p_{2})=F(p_{2})$. Suppose $F(p_{1})\cap F(p_{2})\neq\emptyset$. Then there exist a finite ordered metric space $D^{\prime}$ extending $D$, order-preserving partial isometries $q_{1}$ and $q_{2}$ of $D^{\prime}$ extending $p_{1},p_{2}$ respectively, and an element $w\in\mathsf{F}(s,t)$, the free group generated by s,t{s,t}, such that
@@
Intuitively, $g_{1}$ sends $w(y)$ to a copy which is smaller than $b_{n}$ and fixes $A$ pointwise. Consequently, we have $w_{1}(D)<b_{n}$ and, for any $y\in D$, $w_{1}(y)<x_{0}$ if and only if $w(y)<x_{0}$.
@@
Define an order-preserving partial isometry $g_{2}\colon B\cup w_{1}(D)\to B\cup w_{2}(D)$ by letting $g_{2}(b)=b$ for all $b\in B$ and $g_{2}(w_{1}(y))=w_{2}(y)$ for all $y\in D$. Moreover, $w_{2}(D)<a_{n}$ and, for any $y\in D$, $w_{2}(y)<x_{0}$ if and only if $w_{1}(y)<x_{0}$.
@@
[…]
@@
Toward a contradiction, assume $V\cdot x$ is infinite for all $x\in\mathbb{U}_{\Delta}^{<}$. We prove that $V$ is dense. Since $V$ is also closed, it follows that $V={\rm Aut}(\mathbb{U}_{\Delta}^{<})$, contradicting the properness of $V$.
[…]
@@
[…]
@@
[…]